Rachel invested her savings in two investment funds. The $5000 that she invested in Fund A returned a 3% profit. The amount that she invested in Fund B returned a 10% profit. How much did she invest in Fund B, if both funds together returned a 5% profit?
Answer by mananth(16946)

Fund A ---$ 5000---------3% Interest = $150
Fund B ---$x-------------10% Interest = 0.1x
Total amount invested = 5000+x
Rate of return = 5%
Interest on total amount = 0.05(5000+x)
Fund A interest + Fund B interest = Total Interest
150+0.1x=0.05(5000+x)
150+0.1x= 250+0.05x
0.1x-0.05x= 250-150
0.05x=100
x= 100/0.05
x=2000
$ 2000 was invested in Fund B